OpenStudy (anonymous):

@ganeshie8 What is the length of a side of a cube with a volume of 729 m3?

OpenStudy (anonymous):

HELP

ganeshie8 (ganeshie8):

hint : Volume = \(\large s^3\)

ganeshie8 (ganeshie8):

\(\large s^3 = 729\) solve \(\large s\)

OpenStudy (anonymous):

take the cube root of 729

OpenStudy (anonymous):

9

ganeshie8 (ganeshie8):

Yep !

OpenStudy (anonymous):

What is the side length of a cube with a volume of 2744 mm3?

ganeshie8 (ganeshie8):

do the same thing

OpenStudy (anonymous):

14

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

What is the volume of a cube, if the length of one side is 7 feet? A. 343 cubic feet B. 21 cubic feet C. 49 cubic feet D. 294 cubic feet

ganeshie8 (ganeshie8):

use the same formula (look at hint)

OpenStudy (anonymous):

C

ganeshie8 (ganeshie8):

no u need to cube

ganeshie8 (ganeshie8):

Volime = s^3

OpenStudy (anonymous):

A

ganeshie8 (ganeshie8):

yes!

OpenStudy (anonymous):

A closet is 6 feet long, 3 feet wide, and 7 feet high. What is the volume of the closet?

OpenStudy (anonymous):

6x3x7=?

OpenStudy (anonymous):

126

ganeshie8 (ganeshie8):

correct !

OpenStudy (anonymous):

@ganeshie8 The surface area of a cube is 864 in2. What is the area of one face of this cube?

ganeshie8 (ganeshie8):

how many faces a cube has ?

OpenStudy (anonymous):

12

ganeshie8 (ganeshie8):

nope, try again.

ganeshie8 (ganeshie8):

a cube is just a box

ganeshie8 (ganeshie8):

it has 6 faces : 1) top 2) bottom 3) left 4) right 5) front 6) back

ganeshie8 (ganeshie8):

so volume of one face = 864/6 = ?

OpenStudy (anonymous):

144